③ Let 2,= 2-3 , z = 1+ 2i Find (22) 2 0 ③ compute a Ct - ;) e ① & 2 + 1 - ② i + ti
Answer #1

إلي وه 3 ( 9 +1- رنو - )= (و ) 0 ا-3ر6 + 4 + 3 + في - 13 + و - (2) را -۱) وه لم = ر - راع - 1) و 4- [ ) ar ) + (د) يا 4 - [ 4